log6(x+8)=log6 (x-4)+ 1/3log6 8
Answer by edjones(7569)

log6(x+8)=log6 (x-4)+ 1/3log6 (8)
x+8=(x-4)*2 [8^(1/3)=2]
x+8=2x-8
x=16
.
Ed